OpenStudy (ariellesircy):

Solving Quadratic Equations by Factoring I'm stumped on this one... could somebody explain it to me??? n^2 - 10n + 22 =-2

OpenStudy (jdoe0001):

have you done quadratic factoring yet?

OpenStudy (ariellesircy):

Yes. I am just stuck on the part of I know you have to use 2 numbers that times to get the 10n but what do i do with the 22?

OpenStudy (jdoe0001):

\(\large n^2 - 10n + 22 =-2\implies \begin{array}{cccllll} n^2 &- 10n &+ 24 =0\\ &\uparrow &\uparrow\quad \\ &-6-4&-6\cdot -4 \end{array} \)

OpenStudy (ariellesircy):

Omg... I feel really blonde now because I already knew that... Sorry to bother you.!

OpenStudy (ariellesircy):

Can I ask one more thing??? How would you set it up though. Like n(-6-4) and(or) n(-6x-4)???

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} n^2 &- 10n &+ 24 =0\\ &\uparrow &\uparrow\quad \\ &-6-4&-6\cdot -4 \end{array} \implies (x-6)(x-4)=0\)

OpenStudy (jdoe0001):

ahemm is "n" =) so \(\large \begin{array}{cccllll} n^2 &- 10n &+ 24 =0\\ &\uparrow &\uparrow\quad \\ &-6-4&-6\cdot -4 \end{array} \implies (n-6)(n-4)=0\)

OpenStudy (ariellesircy):

Okayy Thankyou!!!

OpenStudy (jdoe0001):

yw